A single phase overhead transmission line delivers 4000 kW at 11 kV at 0·8 p.f. lagging. If resistance and reactance per conductor are 0·15 Ω and 0·02 Ω respectively, calculate:

(i) percentage regulation

(ii) sending end power factor

(iii) line losses


1
Expert's answer
2021-01-19T05:19:58-0500

"P = 4KW, V_R=11KV, cos(\phi_R)=0.8, r=0.15\Omega, X=0.02\Omega"


"I = \frac{P}{V_R cos(\phi_R)} = \frac{4000\times 100}{11000\times 0.8}=454.54A"


(i) Percentage regulation, "= \frac{V_s-V_R}{V_R}\times 100 = \frac{V_R+I(r+jX)-V_R}{V_R}\times 100 = \frac{454.54(0.15+0.02j)}{11000} \times 100 = 0.624"


(ii) Sending end power factor, "\frac{V_R cos(\phi_R) +IR}{V_s} = \frac{(11000\times 0.8)+(454.54\times 0.15)}{(11000\times 0.8)+(454.54\times (0.15+0.02j))}=0.99"



(iii) Line loss, "I^2R = (454.54^2 \times 0.15) = 30.98 5 KW"